OpenStudy (anonymous):

Solve the equation. Show your work. 16^3 = 4^x

OpenStudy (anonymous):

write it out as exponentials with the same base: 16^3 = 4^x (2^4)^3 = (2^2)^x 2^12 = 2^(2x) since the bases are the same, the exponents must be the same: 12 = 2x can you solve it from here?

OpenStudy (anonymous):

x=6 Thank you

OpenStudy (anonymous):

yw